# Privacy-Preserving Retrieval Augmented Generation with Differential Privacy

### v1

@@ A.1 Main Results across Models and Datasets
$\epsilon_{\mathrm{token}}=1$ 0.04 0.12 0.16 0.04 0.12 0.18 0.04 0.12 0.18 0.04 0.12 0.18
$\epsilon_{\mathrm{token}}=2$ 0.18 0.19 0.19 0.27 0.34 0.33 0.31 0.38 0.39 0.31 0.38 0.39
$\epsilon_{\mathrm{token}}=5$ 0.05 0.06 0.06 0.23 0.22 0.22 0.36 0.40 0.37 0.45 0.47 0.45
A.2 Further Results on Trivia Dataset with OPT (1.3B)
@@
Figure 5: Average match accuracy for questions with different numbers of ground truth relevant documents across baseline algorithms and our algorithms with OPT (1.3B) on Trivia dataset.
Figure 5 shows the performances for different numbers of ground truth relevant documents. We see questions with more relevant documents tend to be answered correctly across algorithms, even with Non-RAG. We also observe that the benefit of having more voters (larger $m$) is larger for small $\epsilon_{\mathrm{token}}$. Interestingly, the benefit is significant when there are less relevant documents for DPVoteRAG and when there are more relevant documents for DPSparseVoteRAG. We don’t see significant difference between $m=30$ and $50$ when $\epsilon_{\mathrm{token}}=5$.
@@
Figure 6: Average numbers of generated tokens for each $\epsilon_{\mathrm{token}}$ and $\epsilon_{\mathrm{total}}$ with OPT (1.3B) on Trivia dataset. We fix $m=50$.
@@
Voting Frequency of DPSparseVoteRAG.
Figure 7: Histograms of voting frequency of DPSparseVoteRAG for each $\epsilon_{\mathrm{token}}$ with OPT (1.3B) on Trivia dataset. We fix $\epsilon_{\mathrm{total}}=20$ and $m=50$.
Figure 7 demonstrates how often DPSparseVoteRAG carries out the voting procedure. We see that DPSparseVoteRAG uses the voting procedure only for the 10% of tokens in the total output for most of the questions when $\epsilon_{\mathrm{token}}\geq 2$. On the other hand, when $\epsilon_{\mathrm{token}}=1$, it uses the voting more often. The results validate that the sparse vector technique helps spend a privacy budget in a principled manner. Furthermore, the results imply that small $\epsilon_{\mathrm{token}}$ can lead to the unnecessary use of the voting, degrading the overall performance.

### v2

@@ B.1 Main Results across Models and Datasets
$\epsilon_{\mathrm{token}}=1$ $0.38$ $0.55$ $0.65$ $0.38$ $0.55$ $0.66$ $0.38$ $0.55$ $0.66$ $0.38$ $0.55$ $0.66$
$\epsilon_{\mathrm{token}}=2$ $0.64$ $\bm{0.72}$ $\bm{0.72}$ $0.71$ $0.78$ $\bm{0.80}$ $0.71$ $0.79$ $\bm{0.83}$ $0.71$ $0.79$ $0.83$
$\epsilon_{\mathrm{token}}=5$ $0.48$ $0.49$ $0.48$ $0.74$ $0.75$ $0.74$ $0.79$ $0.80$ $0.79$ $0.85$ $\bm{0.87}$ $0.85$
Table 9: Average match accuracy comparison of our algorithms on Trivia dataset with Llama 3.2 (1B) under varying values of total privacy budget $\epsilon_{\mathrm{total}}$ with different hyperparameters, the number of voters $m$ and $\epsilon_{\mathrm{token}}$. Bold font values represent the best performance of each algorithm under fixed $\epsilon_{\mathrm{total}}$. We report the means of average match accuracy over three runs.
$m$ 30 40 50 30 40 50 30 40 50 30 40 50
DPVoteRAG
$\epsilon_{\mathrm{token}}=1$ $0.46$ $0.49$ $\bm{0.52}$ $0.54$ $0.57$ $\bm{0.60}$ $0.54$ $0.57$ $0.60$ $0.54$ $0.57$ $0.60$
$\epsilon_{\mathrm{token}}=2$ $0.27$ $0.27$ $0.28$ $0.51$ $0.52$ $0.52$ $0.60$ $\bm{0.61}$ $0.60$ $0.60$ $\bm{0.61}$ $0.60$
$\epsilon_{\mathrm{token}}=5$ $0.10$ $0.10$ $0.10$ $0.27$ $0.27$ $0.28$ $0.47$ $0.47$ $0.47$ $0.56$ $0.57$ $0.56$
DPSparseVoteRAG
$\epsilon_{\mathrm{token}}=1$ $0.38$ $0.52$ $0.63$ $0.38$ $0.52$ $0.63$ $0.38$ $0.52$ $0.63$ $0.38$ $0.52$ $0.63$
$\epsilon_{\mathrm{token}}=2$ $0.61$ $0.62$ $\bm{0.64}$ $0.70$ $0.76$ $\bm{0.77}$ $0.71$ $0.78$ $\bm{0.79}$ $0.71$ $0.78$ $0.79$
$\epsilon_{\mathrm{token}}=5$ $0.43$ $0.43$ $0.45$ $0.66$ $0.64$ $0.66$ $0.76$ $0.75$ $0.76$ $\bm{0.81}$ $0.80$ $0.79$
Table 10: Average match accuracy comparison of our algorithms on NQ dataset with Llama 3.2 (1B) under varying values of total privacy budget $\epsilon_{\mathrm{total}}$ with different hyperparameters, the number of voters $m$ and $\epsilon_{\mathrm{token}}$. Bold font values represent the best performance of each algorithm under fixed $\epsilon_{\mathrm{total}}$. We report the means of average match accuracy over three runs.
$m$ 30 40 50 30 40 50 30 40 50 30 40 50
DPVoteRAG
$\epsilon_{\mathrm{token}}=1$ $0.24$ $\bm{0.27}$ $\bm{0.27}$ $0.31$ $0.38$ $\bm{0.39}$ $0.31$ $0.38$ $0.39$ $0.31$ $0.38$ $0.39$
$\epsilon_{\mathrm{token}}=2$ $0.15$ $0.15$ $0.12$ $0.30$ $0.31$ $0.27$ $0.42$ $\bm{0.44}$ $0.42$ $0.42$ $\bm{0.44}$ $0.42$
$\epsilon_{\mathrm{token}}=5$ $0.05$ $0.05$ $0.05$ $0.15$ $0.15$ $0.13$ $0.30$ $0.30$ $0.27$ $0.38$ $0.39$ $0.35$
DPSparseVoteRAG
$\epsilon_{\mathrm{token}}=1$ $0.14$ $0.25$ $0.32$ $0.14$ $0.25$ $0.34$ $0.14$ $0.25$ $0.34$ $0.14$ $0.25$ $0.34$
$\epsilon_{\mathrm{token}}=2$ $0.31$ $\bm{0.33}$ $\bm{0.33}$ $0.41$ $0.49$ $\bm{0.52}$ $0.42$ $0.50$ $\bm{0.53}$ $0.42$ $0.50$ $0.53$
$\epsilon_{\mathrm{token}}=5$ $0.15$ $0.14$ $0.13$ $0.40$ $0.36$ $0.32$ $0.53$ $0.51$ $\bm{0.53}$ $0.57$ $0.56$ $\bm{0.58}$
Table 11: Average match accuracy comparison of our algorithms on Trivia dataset with GPT2-XL under varying values of total privacy budget $\epsilon_{\mathrm{total}}$ with different hyperparameters, the number of voters $m$ and $\epsilon_{\mathrm{token}}$. Bold font values represent the best performance of each algorithm under fixed $\epsilon_{\mathrm{total}}$. We report the means of average match accuracy over three runs.
$m$ 30 40 50 30 40 50 30 40 50 30 40 50
DPVoteRAG
$\epsilon_{\mathrm{token}}=1$ $0.31$ $0.32$ $\bm{0.34}$ $0.36$ $0.38$ $\bm{0.42}$ $0.36$ $0.38$ $0.42$ $0.36$ $0.38$ $0.42$
$\epsilon_{\mathrm{token}}=2$ $0.22$ $0.23$ $0.23$ $0.37$ $0.38$ $0.37$ $0.44$ $\bm{0.46}$ $\bm{0.46}$ $0.44$ $\bm{0.46}$ $\bm{0.46}$
$\epsilon_{\mathrm{token}}=5$ $0.08$ $0.08$ $0.08$ $0.23$ $0.23$ $0.23$ $0.34$ $0.34$ $0.33$ $0.44$ $0.43$ $0.42$
DPSparseVoteRAG
$\epsilon_{\mathrm{token}}=1$ $0.25$ $0.30$ $0.38$ $0.25$ $0.30$ $0.38$ $0.25$ $0.30$ $0.38$ $0.25$ $0.30$ $0.38$
$\epsilon_{\mathrm{token}}=2$ $0.37$ $\bm{0.39}$ $0.38$ $0.41$ $0.46$ $\bm{0.47}$ $0.42$ $0.47$ $0.48$ $0.42$ $0.47$ $0.48$
$\epsilon_{\mathrm{token}}=5$ $0.28$ $0.27$ $0.27$ $0.43$ $0.43$ $0.43$ $0.52$ $\bm{0.53}$ $\bm{0.53}$ $\bm{0.59}$ $0.57$ $0.56$
Table 12: Average match accuracy comparison of our algorithms on NQ dataset with GPT2-XL under varying values of total privacy budget $\epsilon_{\mathrm{total}}$ with different hyperparameters, the number of voters $m$ and $\epsilon_{\mathrm{token}}$. Bold font values represent the best performance of each algorithm under fixed $\epsilon_{\mathrm{total}}$. We report the means of average match accuracy over three runs.
$m$ 30 40 50 30 40 50 30 40 50 30 40 50
DPVoteRAG
$\epsilon_{\mathrm{token}}=1$ $0.21$ $0.24$ $\bm{0.27}$ $0.25$ $0.28$ $\bm{0.33}$ $0.25$ $0.28$ $0.33$ $0.25$ $0.28$ $0.33$
$\epsilon_{\mathrm{token}}=2$ $0.18$ $0.19$ $0.19$ $0.30$ $0.29$ $0.29$ $0.35$ $0.35$ $\bm{0.36}$ $0.35$ $0.35$ $\bm{0.36}$
$\epsilon_{\mathrm{token}}=5$ $0.08$ $0.08$ $0.07$ $0.19$ $0.19$ $0.19$ $0.30$ $0.29$ $0.29$ $\bm{0.36}$ $0.34$ $0.34$
DPSparseVoteRAG
$\epsilon_{\mathrm{token}}=1$ $0.14$ $0.19$ $0.24$ $0.14$ $0.19$ $0.24$ $0.14$ $0.19$ $0.24$ $0.14$ $0.19$ $0.24$
$\epsilon_{\mathrm{token}}=2$ $0.24$ $0.28$ $\bm{0.31}$ $0.28$ $0.33$ $\bm{0.36}$ $0.29$ $0.35$ $0.40$ $0.29$ $0.35$ $0.40$
$\epsilon_{\mathrm{token}}=5$ $0.18$ $0.18$ $0.16$ $0.34$ $0.33$ $0.35$ $\bm{0.41}$ $0.39$ $\bm{0.41}$ $0.42$ $0.41$ $\bm{0.43}$
B.2 Further Results on Trivia Dataset with OPT (1.3B)
@@
Figure 7: Average match accuracy for questions with different numbers of ground truth relevant documents across baseline algorithms and our algorithms with OPT (1.3B) on Trivia dataset. For VoteRAG, DPVoteRAG, and DPSparseVoteRAG, we report the means and standard deviations over three runs.
Figure 7 shows the performances for different numbers of ground truth relevant documents. We see questions with more relevant documents tend to be answered correctly across algorithms, even with Non-RAG. We also observe that the benefit of having more voters in DPSparseVoteRAG (larger $m$) is larger for small $\epsilon_{\mathrm{token}}$. We don’t see significant difference between $m=30$ and $50$ when $\epsilon_{\mathrm{token}}=5$.
@@
Figure 8: Average numbers of generated tokens for each $\epsilon_{\mathrm{token}}$ and $\epsilon_{\mathrm{total}}$ with OPT (1.3B) on Trivia dataset. We fix $m=50$. We report the means and standard deviations over three runs.